# Enhancing Quantum Memory Lifetime with Measurement-Free Local Error Correction and Reinforcement Learning

### v1

@@ B.4 Details on environment
Data: Code type $T$; lattice size $L$; ambient error rate $p_{amb}$; gate error rate $p_{gate}$; sample size $N$
Function ComputeReward(circuit, $T$, $L$, $p_{amb}$, $p_{gate}$, $N$):
@@
                   data := GenerateError(data, $p_{amb}$, $T$);
@@
Function GenerateError(qubits, $p_{err}$):
@@
Also, we choose specific parameters of the RL environment in each code as shown in Table 1. Note that these parameters were chosen based on comparing the final rewards from training with various choices of this parameter set. In particular, we choose $R$ to be as small as possible for each code to minimize the time for reward computation. However, for 2D toric code, we observe that $R\leq 4$ causes the occurrence of repeating syndrome extraction, which introduces more gate errors into the system without benefiting decoding. This is because, for such a small $R$, the RL agent is not penalized enough by a bad reward to avoid repeating syndrome extraction and to choose the “do nothing” ($\emptyset$) action instead. This is why we choose $R=5$ for the 2D toric code case.
Appendix C Error Bar Computation with Data
We discuss how each of these error bars for varying parameters in this paper is computed as below:
  • Final reward (in Fig. 7) and final logical error rate $P_{L}$ (in Fig. 12):
@@ Appendix C Error Bar Computation with Data
    Recall that the final reward is either a decoding success rate (for 2D and 4D toric code) or a “survival” rate of an initial spin direction (for the 2D Ising model). Note that the final logical error rate $P_{L}$ is defined as a decoding failure rate for 2D toric code. Thus, we consider the final reward (or logical error rate) $\hat{p}$ computed from large $N$ samples in each code as a mean of some binomial distribution, i.e. the estimated final reward is given by the 95% confidence interval with
    $\hat{p}\pm 1.96\times\sqrt{\frac{\hat{p}(1-\hat{p})}{N}}.$ (15)
@@
    We simulate a memory lifetime $T_{i}$ for 1000 samples. Then, we compute the 95% confidence interval for an estimate for average lifetime $T$ as
    $\text{Mean}_{i}(T_{i})\pm 1.96\times\frac{\text{Stdev}_{i}(T_{i})}{\sqrt{1000}}.$ (16)
  • Effective code distance $D_{eff}$ (in Fig. 11) and another fitting parameter $k_{1}$ (in Fig. 15):
@@
    Using the symfit library with setting absolute_sigma = False [90], we fit $D_{eff}$ and other parameters and obtain a value $\hat{D}_{eff}$ and a standard deviation $\sigma_{D_{eff}}$ for each parameter. Then, we compute the 95% confidence interval for $D_{eff}$ as
    $\hat{D}_{eff}\pm 1.96\times\sigma_{D_{eff}}.$ (17)
    We did the same for $k_{1}$ as well.
Based on these statistical methods, we obtained the errorbar in Fig. 7, 10, 11, 12, and 15.
Appendix D Classifying remaining errors after LEC
Figure 13: Classification of remaining errors after Toom LEC circuit (grey) and RL-optimized LEC circuit (blue) in each code with 10,000 samples. Each sample is classified into one of the categories labeled on the x-axis. Note that this statistic is based on 10,000 samples with the code and error model parameters as in Table 1 with $p_{gate}=1\times 10^{-3}$ for 2D Ising model and $p_{gate}=1\times 10^{-5}$ for 4D toric code. By categorizing this data more coarsely, we obtain the classification as shown in Fig. 8. (a) Let the number of flipped spins be $N_{F}$. Then, $(a,b)$ implies that $a<N_{F}<b$, and $[a,b]$ implies that $a\leq N_{F}\leq b$. (b) Distribution for flipped spins between 8 and 56. Note that “spikes” are formed at the multiples of $L$. (c) The label wider sheet refers to the error sheets wider than $(1,1)$, and the label uncorrect. etc. refers to the uncorrectable error patterns that are not categorized as the presence of a single error sheet.
For the classification of error patterns as shown in Fig. 8, we obtain 10,000 samples, which are error-correcting code lattices with multiple LEC rounds with independent ambient and gate errors being introduced, and count the number of each error pattern. For 2D toric code, we classify the error patterns based on individual length-$N$ error chains within 10,000 lattices after 5 LEC rounds. Thus, each lattice can have multiple length-$N$ error chains. On the other hand, for self-correcting memories, we classify the error patterns based on the total number of flipped spins (for 2D Ising) or the presence of width-$N$ error sheets (for 4D toric). Thus, each lattice falls into one of the categories. This is why in Fig. 8, 2D toric code has more counts than self-correcting memories.
The single LEC action in the labels of Fig. 8 refers to the error removal operations that are included as the candidate operations as discussed in Appendix A. Note that these operations consist of the RL-optimized LEC circuit but not conventional LEC circuits.
For the 2D toric code, we label each length-$N$ error chain as a non-logical error correctable by single LEC action if $N\leq 3$. Otherwise, we label it as a non-logical error uncorrectable by single LEC action if it is not a logical error.
For the 2D Ising model, we first classify each sample lattice based on the number of flipped spins among $648^{2}=$ 64 data qubits as shown in Fig. 13(a). From Fig. 13(b), we already observe that an RL-optimized LEC circuit is more effective in suppressing the occurrence of a non-logical error uncorrectable by a single LEC action compared to the Toom LEC circuit. This is consistent with the analysis in Section V.1. To obtain a simplified classification as in Fig. 8, we perform a perfect Toom’s rule decoding — just as for 4D toric code. Then, we label a sample lattice as a non-logical error correctable by single LEC action if the remaining error pattern is either a width-1 or width-2 error sheet, a non-logical error uncorrectable by single LEC action if the remaining error pattern is width-$N$ (where $83\leq N<$ 8) error sheet, and a logical error undetectable by syndromes if the remaining error pattern is a width-8 error sheet.
For the 4D toric code, we classify each sample lattice based on error patterns after 2 LEC rounds followed by a perfect Toom’s rule decoding. Note that we observe some uncorrectable error patterns that do not fall into the category of width-$N$ error sheets as shown in Fig. 13(c). We label both samples with wider sheets and such miscellaneous uncorrectable errors as non-logical error uncorrectable by single LEC action. Just as in Section V.1, we also observe that an RL-optimized LEC circuit is more effective in suppressing the occurrence of not only the wider error sheets but also such miscellaneous uncorrectable errors.
Appendix E Lifetime vs. size for 2D toric code
@@ Appendix D Classifying remaining errors after LEC
As shown in Fig. 10 of Section V.3, we observe that $T$ gets slightly decreased as $L$ gets increased for the RL-optimized LEC circuit. After some investigation, we conclude that this is because length-$N$ error chains contribute differently to $T$. As $L$ gets larger, final recovery by perfect syndrome extraction with perfect ancilla readout and MWPM classical algorithm can correct longer error chains. Thus, the effect of LEC circuits in extending the memory lifetime gets weaker due to the varying power of final global decoding as $L$ varies.
@@
Figure 14: Comparison between data with varying $L$ benchmarked with perfect gates. (a) Average memory lifetime $T$ vs. benchmark lattice size $L$ for an LEC circuit trained with $L=8$ (blue circle) and an LEC circuit trained with $L=12$ (orange square). For both training, $H_{\max}=40$, $p_{amb}=0.02$, $N=100$, $R=7$, and $p_{gate,train}=1\times 10^{-4}$. (b) Number of error chains out of 1 million samples divided by $L^{2}$ for each category of error chain after two LEC rounds. For each benchmark $L$ case, we use the common RL-optimized LEC circuit trained at $L=8$, $H_{\max}=40$, $p_{amb}=0.02$, $N=100$, $R=7$, and $p_{gate,train}=1\times 10^{-4}$. Note that logical error occurs in only three samples for the $L=8$ case and none for other $L$s.
This explanation is supported by Fig. 14:
  • In Fig. 14(a), we find that $T$ decreases as benchmark $L$ increases for both RL-optimized LEC circuits with perfect gates trained at $L=8$ and $L=12$, respectively. Especially, we observe the “drop” of lifetime between benchmark $L=8$ and $L=12$. This implies that the relation between a lifetime and lattice size is independent of the training lattice size for optimizing the LEC circuits.
@@
  • In Fig. 14(b), we observe that each length-$N$ error chains occur with a similar density regardless of benchmark $L$ for a fixed LEC circuit with perfect gates. This implies that before the final recovery, the effect of the LEC circuit in reducing the number of error chains is similar regardless of benchmark $L$. Thus, we explain this relation between a lifetime and lattice size based on the role of final recovery.
@@
Appendix F Other parameters from log-log fitting between lifetime and ambient error rate
Although Section V.3 focuses on the fitting parameter $D_{eff}$ from fitting $\log T$ and $\log(p_{amb})$, there exist other fitting parameters as well: $k$ for 2D toric code and $k_{1}$ and $k_{2}$ for self-correcting memories. By comparing Equation (12) and (13) with (11), we observe that
$k$ $=\log C-D_{eff}\times\log\left(\alpha\right)$ (18)
$k_{1}$ $=\log\left(\alpha\right)$ (19)
$k_{2}$ $=\log C$ (20)
@@ Appendix F Other parameters from log-log fitting between lifetime and ambient error rate
For 2D Ising model and 4D toric code, we plot the fitted $k_{1}$ vs. benchmark $p_{gate}$ as shown in Fig. 15 to check whether its behavior is consistent with the intuition on $\alpha$ discussed in Section V.2. Note that we also confirm from the fitting that the standard deviation for $k$ and $k_{2}$ do not diverge as well.
@@
Figure 15: $k_{1}$ vs. Benchmark $p_{gate}$ for RL-optimized LEC circuits trained at $p_{gate,1}$ (blue circle) and $p_{gate,2}$ (orange diamond) as well as Toom LEC circuit (grey square) in (a) 2D Ising model and (b) 4D toric code. Our computation for the error bar of $k_{1}$ is discussed in Appendix C.
@@
Note that $k_{1}$ shows different values for different $p_{gate}$ and LEC circuits. We expect that $\alpha$ varies for different $p_{gate}$ and LEC circuits. From Equation (19), increasing $k_{1}$ implies increasing $\alpha$, because a logarithm function is monotonically increasing. Recall that in Section V.2, we discuss that $\alpha$ denotes how many new errors get introduced every LEC round. Then, we can explain two general trends observed in Fig. 15 based on this intuition on $\alpha$:
  • For each LEC circuit, $k_{1}$ increases as benchmark $p_{gate}$ increases. This trend is consistent with the intuition on $\alpha$ — as $p_{gate}$ increases, more new errors are introduced to the system.
@@
  • For each benchmark $p_{gate}$, $k_{1}$ is the biggest for the RL-optimized LEC circuit trained with smaller $p_{gate}$ and the smallest for the Toom LEC circuit. Note that in Section IV.2, we show that RL-optimized LEC circuit trained with larger $p_{gate}$ includes less higher-weight error removal operations, such as $d=1$ and $d=2$ actions for 2D Ising model or $d=1$ and $d=(1,1)$ actions for 4D toric code. We analyze that this is because the higher-weight error removal operations may introduce more errors into the system compared to the Toom error removal operation. Especially, for 4D toric code, we “grouped” parallelized $d=1$ and $d=(1,1)$ operations each into single actions (see Appendix A). Thus, for the same benchmark $p_{gate}$, the most “complicated” LEC circuits that are trained with smaller $p_{gate}$ show the biggest $k_{1}$, and the “simplest” Toom circuits show the smallest $k_{1}$.
Appendix G LEC circuits with large gate error rate
For 2D toric code, RL training with large $p_{gate}$ results in the simplest NN decoder. Note that we fix our first action in the LEC circuit to be a syndrome extraction (see Appendix B.2). Thus, we do not observe that the RL framework produces a “do nothing” LEC circuit. However, we find that for $p_{gate}>0.01$, the simplest NN circuit performs worse than just “do nothing.”
@@ Appendix G LEC circuits with large gate error rate
Figure 16: $D_{eff}$ vs. Benchmark $p_{gate}$ for RL-optimized LEC circuit trained at large $p_{gate,train}$ (green circle) as well as Toom LEC circuit (grey square) in (a) 2D Ising model and (b) 4D toric code. Here, we choose $p_{gate,train}$ to be $9\times 10^{-3}$ for 2D Ising model and $3\times 10^{-5}$ for 4D toric code.

### v2

@@ C.2 Details on environment
Data: Code type $T$; lattice size $L$; ambient error rate $p_{\rm{amb}}$; gate error rate $p_{\rm{gate}}$; sample size $N$
Function ComputeReward(circuit, $T$, $L$, $p_{\rm{amb}}$, $p_{\rm{gate}}$, $N$):
@@
                   data := GenerateError(data, $p_{\rm{amb}}$, $T$);
@@
Function GenerateError(qubits, $p_{\rm{err}}$):
@@
After multiple LEC rounds, we evaluate the decoding success. This final recovery step is to determine whether the final state of data qubits after multiple LEC rounds is “close enough” to the trivial ground state of the error-correcting code. This step varies among the choice of the code. For 2D toric code, we perform global decoding by minimum-weight perfect matching (MWPM) classical algorithm with precise syndrome information using PyMatching package [14, 16]. For 2D Ising model, we perform a majority vote among data qubits. For 4D toric code, we perform enough (50 rounds) repetition of perfect Toom’s rule decoding. After this step, 2D and 4D toric code decoding are evaluated to be successful if we get a trivial ground state, whereas 2D Ising model decoding is evaluated to be successful if we get less than 50% of all spins bit-flipped. We repeat this simulation for $N$ times with a fixed LEC circuit. For each simulation, a stochastic error model results in different configurations of errors applied to qubits. Thus, we can obtain the decoding success rate through this benchmark simulation.
As discussed in Section III.2, the uncorrectable error configurations after perfect Toom’s rule decoding in 4D toric code can have a low error density. In principle, these uncorrectable errors should be decoded by some non-local decoding protocol [44]. However, we decide to leave these uncorrectable errors as decoding failures, because (1) the non-local decoding protocol takes significant running time and (2) the prior LEC reduces the uncorrectable error configurations with low error density.
For 2D and 4D toric code, the reward is computed as the number of successful decoding cases divided by the number of total cases $N$. However, for the 2D Ising model, the reward is computed as an average non-error density, i.e. average number of non-flipped spins divided by the total number of spins $L^{2}$.
@@
In particular, we choose $R$ in Table 4 to be as small as possible for each code to minimize the time for reward computation. However, for 2D toric code, we observe that $R\leq 4$ causes the occurrence of repeating syndrome extraction, which introduces more gate errors into the system without benefiting decoding. This is because, for such a small $R$, the RL agent is not penalized enough by a bad reward to avoid repeating syndrome extraction and to choose the “do nothing” ($\emptyset$) action instead. Thus, we choose $R=5$ for the 2D toric code case.
C.3 Conditions on training termination
Each epoch of our RL training consists of 500 episodes with a mini-batch size of 50. In other words, we collect 500 samples of LEC circuits and their benchmarked decoding success rates, divide them into smaller mini-batch subsets with 50 samples each, and perform 10 updates with these subsets on the policy and value network. Before saving our updated neural network model, we run 40 epochs for the 2D toric code, 80 epochs for the 2D Ising model, and 40 epochs for the 4D toric code. We terminate each RL training when a neural network model produces the same LEC circuit as output as the previously saved one throughout these epochs. For 2D toric code with $H_{\max}=40$ and $p_{\rm{gate}}=1\times 10^{-4}$, each RL training requires 200-400 epochs. For self-correcting memories with $H=60$ and any $p_{\rm{gate}}$, each RL training requires 800-1000 epochs for the 2D Ising model and 400-600 epochs for the 4D toric code. Note that these are obtained with a fully connected neural network with two hidden layers of size 128. Also, we use a variable depth agent for 2D toric code and a fixed depth agent for 2D Ising model and 4D toric code (see Appendix C.1).
C.4 Details on RL hyperparameters
As shown in Table 5, we compare the final reward from RL training with a fully connected neural network with two hidden layers of size $d_{\rm{NN}}$ with four independent runs. We observe that the final reward gets improved until $d_{\rm{NN}}=128$. However, for $d_{\rm{NN}}>128$, we observe that a similar reward was obtained with shorter training epochs before termination. Considering the size of saved neural networks after training, we choose $d_{\rm{NN}}=128$ for our RL training. Also, this is why we did not have to implement a more complicated neural network structure — such as a transformer — because a simple fully connected neural network with two hidden layers of size 128 was enough for our optimization task.
C.5 Stability of RL Training
As discussed in Section IV.1, we choose one LEC circuit out of 4 RL executions that maximizes the reward. We record rewards by RL LECs from 4 runs as well as the reward by the conventional LEC with the same parameters as in Table 6. Also, we perform 40 runs of RL training for 2D toric code and 2D Ising model with $p_{\rm{gate}}=1\times 10^{-4}$ and $1\times 10^{-3}$, respectively. As shown in Fig. 12, we record the maximum and minimum rewards by RL LECs out of the first 4 and all 40 runs for each code as well as the reward by the conventional LEC with the same parameters. The final reward of any RL LEC is higher than that of conventional LEC, which we benchmark the decoding performance against. Also, we find that the best final reward out of 4 and 40 runs is marginally different, which justifies the number of RL training executions. These results support the stability of RL training, i.e., RL training results do not vary significantly in performance with respect to our demand.
Figure 12: Final reward comparison of conventional vs. RL LEC chosen from worst or best runs out of multiple training executions. Note that conventional LEC is a nearest-neighbor LEC circuit for 2D toric code and Toom’s rule-based LEC circuit for 2D Ising model and reward is computed out of 10000 independent samples. RL training for 2D toric code and 2D Ising model are done with $p_{\rm{gate}}=1\times 10^{-4}$ and $1\times 10^{-3}$, respectively, and with other parameters in Table 4.
2D Toric Code 2D Ising Model
$(L,H_{\max}N,R,p_{\rm{amb}},p_{\rm{gate}})=(8,40,100,7,0.02,1\times 10^{-4})$ $(L,H_{\max}N,p_{\rm{amb}},p_{\rm{gate}})=(8,60,100,0.40,1\times 10^{-3})$
$d_{\rm{NN}}$ Run 1 Run 2 Run 3 Run 4 $d_{\rm{NN}}$ Run 1 Run 2 Run 3 Run 4
Table 5: RL training results for 2D toric code and 2D Ising model with various network size $d_{\rm{NN}}$. We construct the RL agent with a fully connected neural network with two hidden layers of size $d_{\rm{NN}}$ (for both the policy and value network of a PPO algorithm) and the RL environment with the parameters shown in the table. In each cell, we record a final reward and the required number of epochs before training termination — “Reward (Number of epochs)” — for each $d_{\rm{NN}}$ and training index.
RL LEC - Run 1 RL LEC - Run 2 RL LEC - Run 3 RL LEC - Run 4 Conventional LEC
2D Toric Code 90.62 $\pm$ 0.57 90.29 $\pm$ 0.58 91.46 $\pm$ 0.55 91.54 $\pm$ 0.55 86.39 $\pm$ 0.67
2D Ising Model 90.28 $\pm$ 0.58 90.64 $\pm$ 0.57 89.85 $\pm$ 0.59 89.80 $\pm$ 0.59 86.48 $\pm$ 0.67
4D Toric Code 95.28 $\pm$ 0.42 92.82 $\pm$ 0.51 92.47 $\pm$ 0.52 93.90 $\pm$ 0.47 87.06 $\pm$ 0.66
Table 6: Reward of LEC circuit optimized by 4 RL training executions and conventional LEC circuit. The reward is computed out of 10000 independent samples, and the unit of reward is %. Also, we choose $p_{\rm{gate}}=1\times 10^{-4}$ and $p_{\rm{gate}}=1\times 10^{-5}$ for 2D and 4D toric code, respectively, and $p_{\rm{gate}}=1\times 10^{-3}$ for 2D Ising model (with other parameters as in Table 4). For all training executions in the same code, we use the same code and error model parameters.
This stability of RL training can be explained by a simple state transition, which, in principle, results in low variance in training. In general, RL is applied to problems where the next state $s^{\prime}$ is obtained from the conditional probability distribution $P$, given the current state $s$ and action $a$, i.e., $s^{\prime}\sim P(s^{\prime}|s,a)$ [52, 83]. For example, in the Gym Acrobot game [118], $s^{\prime}$ would be updated from $s$ and $a$ based on stochastic laws of physics, which are hidden from the RL agent and have to be “learned.” However, in our problem, $s^{\prime}$ is determined simply by appending $a$ to $s$. Thus, our RL training is expected to find the optimized solution more reliably compared to the general application of RL methods, which often report substantial training instabilities [119, 120].
Appendix D Error Bar Computation with Data
Each error bar represents a 95% confidence interval for each parameter.
  • Final reward (in Fig. 7) and final logical error rate for 2D toric code $P_{\rm{L}}$ (in Fig. 11):
@@ Appendix D Error Bar Computation with Data
    The final reward is either a decoding success rate (for 2D and 4D toric code) or a “survival” rate of an initial spin direction (for the 2D Ising model). From large $N$ samples noted in the captions, we compute each of their estimates as a mean of a binomial distribution.
@@
    We compute its estimate as a mean of 1000 samples of independently simulated memory lifetime.
  • Effective code distance $D_{\rm{eff}}$ (in Fig. 10) and another fitting parameter $k_{1}$ (in Fig. 15):
@@
    We obtain fitting parameters and their standard deviations with the symfit library [96]. Note that we set the parameter absolute_sigma to be False.
Appendix E Details on memory lifetime extension
E.1 Effect of LEC on error statistics
In Section IV.2, we confirm that the RL LEC performs better than the conventional LEC in each code. More specifically, RL LEC shows a lower decoding failure rate compared to conventional LEC after a certain number of repeated LEC cycles. To analyze the reason for this behavior, we classify remaining error patterns after repeated LEC cycle(s) in each code as shown in Fig. 13. Although both RL and Toom LEC show similar occurrences of logical errors, we observe that RL LEC is effective in suppressing non-logical errors that are either correctable or uncorrectable by a single LEC action that consists of the circuits.
Figure 13: Classification of remaining errors after LEC cycle(s) in each code. Let a single LEC action be any LEC action among extended error removal operations as in Fig. 5. Then, we perform a benchmark simulation for applying LEC cycle(s) with 10000 independent samples. Note that for each LEC circuit, $p_{\rm{gate}}$ is $1\times 10^{-4}$ for 2D toric code, $1\times 10^{-3}$ for 2D Ising model, and $1\times 10^{-5}$ for 4D toric code. Each error pattern after LEC cycle(s) is either (1) a non-logical low-weight error pattern correctable by one of single LEC actions, (2) a non-logical high-weight error pattern uncorrectable by any single LEC action, or (3) a logical error undetectable by extracting syndromes. We compare the occurrence of each category between RL-optimized LEC circuits (blue) vs. conventional LEC circuits in each code (grey).
E.2 Lifetime vs. size for 2D toric code
@@ E.1 Effect of LEC on error statistics
In Fig. 9, we observe that $T$ slightly decreases as $L$ increases for the RL-optimized LEC circuit. We conclude that this is because length-$N$ error chains contribute differently to $T$. As $L$ gets larger, final recovery by perfect syndrome extraction with perfect ancilla readout and MWPM classical algorithm can correct longer error chains. Thus, the effect of LEC circuits in extending the memory lifetime gets weaker due to the varying power of final global decoding as $L$ varies.
@@
Figure 14: Comparison between data with varying $L$ benchmarked with perfect gates. (a) Average memory lifetime $T$ vs. benchmark lattice size $L$ for an LEC circuit trained with $L=8$ (blue circle) and an LEC circuit trained with $L=12$ (orange square). For both training, $H_{\max}=40$, $p_{\rm{amb}}=0.02$, $N=100$, $R=7$, and $p_{\rm{gate,train}}=1\times 10^{-4}$. (b) Number of error chains out of 1 million samples divided by $L^{2}$ for each category of error chain after two LEC rounds. For each benchmark $L$ case, we use the common RL-optimized LEC circuit trained at $L=8$, $H_{\max}=40$, $p_{\rm{amb}}=0.02$, $N=100$, $R=7$, and $p_{\rm{gate,train}}=1\times 10^{-4}$. Note that logical error occurs in only three samples for the $L=8$ case and none for other $L$s.
@@
Fig. 14 supports this explanation. In Fig. 14(a), we find that $T$ decreases as benchmark $L$ increases for both RL-optimized LEC circuits with perfect gates trained at $L=8$ and $L=12$, respectively. Especially, we observe the “drop” of lifetime between benchmark $L=8$ and $L=12$. This implies that the relation between a lifetime and lattice size is independent of the training lattice size for optimizing the LEC circuits. Also, in Fig. 14(b), we observe that each length-$N$ error chains occur with a similar density regardless of benchmark $L$ for a fixed LEC circuit with perfect gates. This implies that before the final recovery, the effect of the LEC circuit in reducing the number of error chains is similar regardless of benchmark $L$. Thus, we explain this relation between a lifetime and lattice size based on the role of final recovery.
@@
E.3 Other parameters from log-log fitting between lifetime and ambient error rate
@@
Although Section V.2 focuses on the fitting parameter $D_{\rm{eff}}$ from fitting $\log_{10}T$ and $\log_{10}(p_{\rm{amb}})$, there exist other fitting parameters as well: $k$ for 2D toric code and $k_{1}$ and $k_{2}$ for self-correcting memories. For 2D Ising model and 4D toric code, we plot the fitted $k_{1}$ vs. benchmark $p_{\rm{gate}}$ as shown in Fig. 15 to check whether its behavior is consistent with the intuition on $\alpha$. Note that we also confirm from the fitting that the standard deviation for $k$ and $k_{2}$ do not diverge as well.
@@
Figure 15: $k_{1}$ vs. Benchmark $p_{\rm{gate}}$ for RL-optimized LEC circuits trained at $p_{\rm{gate},1}$ (blue circle) and $p_{\rm{gate},2}$ (orange diamond) as well as Toom LEC circuit (grey square) in (a) 2D Ising model and (b) 4D toric code. Our computation for the error bar of $k_{1}$ is discussed in Appendix D.
@@
Note that $k_{1}$ shows different values for different $p_{\rm{gate}}$ and LEC circuits. We expect that $\alpha$ varies for different $p_{\rm{gate}}$ and LEC circuits. From Equation (7), increasing $k_{1}$ implies increasing $\alpha$, because a logarithm function is monotonically increasing. Physically, $\alpha$ denotes how many new errors get introduced every LEC round. Then, we can explain two general trends observed in Fig. 15 based on this intuition on $\alpha$.
@@
For each LEC circuit, $k_{1}$ increases as benchmark $p_{\rm{gate}}$ increases. This trend is consistent with the intuition on $\alpha$ — as $p_{\rm{gate}}$ increases, more new errors are introduced to the system. Also, for each benchmark $p_{\rm{gate}}$, $k_{1}$ is the biggest for the RL-optimized LEC circuit trained with smaller $p_{\rm{gate}}$ and the smallest for the Toom LEC circuit. Note that in Section IV.2, we show that RL-optimized LEC circuit trained with larger $p_{\rm{gate}}$ includes less higher-weight error removal operations. This is because the higher-weight error removal operations may introduce more errors into the system compared to the Toom operation. Thus, for the same benchmark $p_{\rm{gate}}$, $k_{1}$ is largest for the most “complicated” LEC circuits that are trained with smaller $p_{\rm{gate}}$ and smallest for the “simplest” Toom circuits.
@@
Figure 16: $D_{\rm{eff}}$ vs. Benchmark $p_{\rm{gate}}$ for RL-optimized LEC circuit trained at large $p_{\rm{gate,train}}$ (green circle) as well as Toom LEC circuit (grey square) in (a) 2D Ising model and (b) 4D toric code. Here, we choose $p_{\rm{gate,train}}$ to be $9\times 10^{-3}$ for 2D Ising model and $3\times 10^{-5}$ for 4D toric code.
E.4 LEC circuits with large gate error rate
For 2D toric code, RL training with large $p_{\rm{gate}}$ results in the simplest NN decoder. Note that we fix our first action in the LEC circuit to be a syndrome extraction (see Appendix C.1). Thus, we do not observe that the RL framework produces a “do nothing” LEC circuit. However, we find that for $p_{\rm{gate}}>0.01$, the simplest NN circuit performs worse than just “do nothing.”